darren and doug must paint the 792 foot fence that encircles thier family home. because darren is older, he has agreed to paint 20% more than darren. how much of the fence will each boy Paint?
Answer by nerdybill(7384)

darren and doug must paint the 792 foot fence that encircles thier family home. because darren is older, he has agreed to paint 20% more than darren. how much of the fence will each boy Paint?
.
Let x = length Doug will paint
then
x + .2x = length Darren will paint
.
x + (x+.2x) = 792
x + x + .2x = 792
2.2x = 792
x = 792/2.2
x = 360 feet (Doug)
.
Darren:
x + .2x = x(1+.2) = 1.2x = 1.2(360) = 432 feet